OpenStudy (anonymous):

Find the sum of all integers between 1900 and 3000, inclusive, which are not divisible by 7.

OpenStudy (anonymous):

I know its going to be multiples of 7 starting from 1904 and ending with 2996

OpenStudy (anonymous):

See, first find all the sum of numbers from 1900 and 3000..

OpenStudy (anonymous):

1900,1901,1902,1903,..........................3000 \(a = 1900\) \(d = 1\) \(l = 3000\) Find number of terms ie \(n\)..

OpenStudy (anonymous):

\[a + (n-1)d = 3000 \\ 1900 + n-1 = 3000 \\ n = 3001 - 1900\]

OpenStudy (anonymous):

What is \(n\) equal to?

OpenStudy (anonymous):

1101

OpenStudy (anonymous):

Good..

OpenStudy (anonymous):

To find sum, use: \[Sum = \frac{n}{2}[a + l]\]

OpenStudy (anonymous):

Find sum now..

OpenStudy (anonymous):

2697450

OpenStudy (anonymous):

Sure?

OpenStudy (anonymous):

Okay, good..!!

OpenStudy (anonymous):

Coming to second part..

OpenStudy (anonymous):

Find sequence now for numbers divisible by 7 in that range..

OpenStudy (anonymous):

1904, 1911, 1918.......................2996..

OpenStudy (anonymous):

\(a = 1904, d = 7, l = 2996\)

OpenStudy (anonymous):

Find \(n\) here..

OpenStudy (anonymous):

Then find the sum and at last, subtract this sum from the sum you found earlier, all easy to do now?

OpenStudy (freckles):

sorry don't mean to interrupt but I think your first answer to @eta 's question has the thousands digit number off

OpenStudy (anonymous):

Which answer @freckles ?

OpenStudy (freckles):

his first answer

OpenStudy (anonymous):

What it is?

OpenStudy (freckles):

the 2697450 answer I have 2694450

OpenStudy (anonymous):

Nope.. He is all good..

OpenStudy (freckles):

no i'm wrong sorry

OpenStudy (anonymous):

\[Sum = \frac{1101}{2} \times 4900\]

OpenStudy (freckles):

i made a type-o in one of my numbers

OpenStudy (anonymous):

It is okay, nothing to worry, at times I am also like you.. :P

OpenStudy (freckles):

Too bad none of us can be gods :p

OpenStudy (anonymous):

There is not need too.. :P

OpenStudy (anonymous):

I think @imyint you can do it now on your own..

OpenStudy (anonymous):

yes thank you so much @eta !!!

OpenStudy (anonymous):

If you come across any doubt, then ask @freckles as I am now going to sleep.. :P

OpenStudy (anonymous):

Okay, good.. \(\dagger\)